Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2007-12-01
ISS On-Orbit Status 12/01/07

Saturday. FE-1 Malenchenko continued preparations for operating the Russian/German TEKh-20 Plasma Crystal-3 Plus (PK-3+) experiment payload.

(After yesterday's hardware setup, leak checking of the electronics box and evacuation of the vacuum work chamber (ZB) with the turbopump, the CDR conducted more testing and calibration, uploaded new software from a USB stick to the payload laptop, checked out the software installation and verified the readiness of the experiment. After additional leak checking on the work chamber during the day, Yuri will deactivate the turbopump tonight at ~4:25pm EST. The experiment is performed on plasma, i.e., fine particles charged and excited by HF (high frequency) radio power inside the evacuated work chamber. Main objective is to obtain a homogeneous plasma dust cloud at various pressures and particle quantities with or without superimposition of an LF (low frequency) harmonic electrical field. The experiment is conducted in automated mode. PK-3+ has more advanced hardware and software than the previously used Russian PKE-Nefedov payload.)

CDR Whitson continued final 10A EVA preparations in the Joint Airlock (A/L) by setting up and starting the periodic scrubbing process on the cooling water loops of EMU (Extravehicular Mobility Unit) #3006 & #3018, afterwards initiating their ionic and particulate filtration, one at a time, with the ion and 3-micron filter. (Purpose of the scrubbing, including iodination of the LCVGs (Liquid Cooling & Ventilation Garments) for biocidal maintenance, is the elimination of any biomass (organic) and particulate matter that may have accumulated in the loops.)

The CDR then terminated the recharging on the first set of EMU helmet light and PGT (Pistol Grip Tool) batteries, started yesterday, and initiated the process on the second set. (The recharged helmet light batteries were then installed in the #3006 & #3018 EMUs, the PGT batteries in PGTs #1001, #1005, #1004 & #1008. Remaining charged batteries were stowed.)

FE-2 Dan Tani (who will return to Earth with STS-122/Atlantis in two weeks) performed the standard inspection and degassing of A/L PWRs (Payload Water Reservoirs, #1012, #1013). (The de-gassing manually removes any air bubbles from the PWR water to minimize the amount of air introduced into the EMU feedwater tanks.)

While waiting for the EMU scrubbing to complete, Peggy Whitson terminated the Fine Leak check conducted overnight on the Lab/Node-2 vestibule, then repressurized the vestibule, restowed the measuring equipment and re-installed the IMV (Intermodular Ventilation) supply-&-return jumpers, two ducts and four O-rings.

The FE-2 meanwhile took two water samples from the EDV water container associated with the hygiene station in the FGB for return on 1E, one from flush water, the other for post-flight microbial analysis.

Peggy & Dan conducted the standard pre-launch SSRMS (Space Station Remote Manipulator System) checkout activities. The SSRMS was then maneuvered by the ground via S-band to the position required for 1E: first a base change (~7:50-8:50am EST) by 'walking off' the Lab PDGF (Power & Data Grapple Fixture), i.e., grappling the Node-2 PDGF and releasing the Lab PDGF, then moving its joints by JOCAS (Joint Operator Commanded Auto Sequence) through two 'waypoints' to the 1E start position.

Yuri Malenchenko performed the semi-annual inspection and photo-documentation of Service Module (SM) windows, tagging up with ground specialists. To check for new defects, photographs of the window panes will be recorded, along with text files, on the RSK1 laptop for subsequent downlink via U.S. OCA assets. (Objective of the inspection, using digital still camera (Nikon D1X w/SB-28DX flash), voice recorder and a flaw log, is to assess the pane surfaces on SM windows 6, 7, 8, 12, 13 for any changes (new cavities, scratches, new or expanded old stains or discolorations affecting transparency properties) since the last inspection, performed by Oleg Kotov on 6/26/07. The new assessment will be compared to the earlier observations.)

Later, Peggy connected the regular ITCS LTL (Internal Thermal Control System/Low Temperature Loop) coolant jumper to the LAB1D6 rack, to support the ground-commanded activation of the U.S. CDRA (Carbon Dioxide Removal Assembly) at ~3:00pm in support of tomorrow's SPHERES experiment by Dan. (The experiment's floating 'satellite' spheres use CO2 as propellant.)

The FE-1 completed handled the daily routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables. (Regular daily SOZh maintenance consists among else of replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.)

In the SM, Malenchenko worked on the ESA/Russian KUBIK-3 refrigerator, setting it up, copying data to laptop, switching the data cable, copying data again, for download via BSR-TM payload telemetry channel, then closed out operations and deactivated the refrigerator. Stowage was to be photographed as a task item from the voluntary 'time permitting' job list.

As a second discretionary task list item, Yuri was to complete another radiation data monitoring & logging session for flow & dose power data with the Matryoshka-R radiation payload and its LULIN-5 electronics box (accumulated readings were recorded on a log sheet for subsequent downlink to TsUP/Moscow via the BSR-TM payload data channel).

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2), TVIS treadmill (FE-1), RED resistive exerciser (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Later, Dan Tani copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

At ~8:50am EST, the crewmembers held their regular weekly planning conference (WPC) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners), via S-band/audio, reviewing upcoming activities and any concerns about future on-orbit events.

At ~9:40am, the FE-1 had his weekly PFC (Private Family Conference) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on the SSC-9 laptop).

Weekly Science Update (Expedition Sixteen -- Weeks 5 & 6)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): From Week 5: The ALTEINO instrument has been rotated and re-activated nominally on 11/22. Subsequent downlink of the files has been performed nominally on 11/23. Radiation measurements continue to be performed throughout the Increment 16 in the PIRS module. From Week 6: Radiation measurements continue nominally in the PIRS module.

ANITA: Completed.

BCAT (Binary Colloidal Alloy Test): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): 'Dan, next week your second and last on-orbit session will be hard scheduled. Thanks for participating and performing your first session as Voluntary Science'.

CFE (Capillary Flow Experiment): Reserve.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): The CSLM SPU (Sample Processing Unit) is currently under vacuum and should be ready to process on Monday.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ETD (Eye Tracking Device): In progress.

Integrated Immune: In progress.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: From Week 5: The EMCS (European Modular Cultivation System) Facility has been switched off on 11/15. MULTIGEN-1 samples will be downloaded on STS-122 (1E).

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.

NUTRITION/REPOSITORY: Dr. Whitson was thanked for completing her Flight Day 30 Nutrition/Repository collections: 'We appreciate the crew notes with all of the barcode and hardware information, they were very helpful'. Her next session is on FD60.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems: Ongoing.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Peggy, we have placed an Actiwatch download/initialization session on the task list for next week. This will include downloading and initializing the lost Actiwatch. Thanks for continuing to sleep log. And thanks also to Dan.'

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress. To be conducted tomorrow (12/2).

Swab (Characterization of Microorganisms & Allergens in Spacecraft): In progress.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

CEO (Crew Earth Observations): Through 11/28 the ground has received a total of 4,017 CEO frames for review and cataloging for Increment 16. 'That includes nearly 1,000 frames in the past week alone! We are focusing our efforts on 14 sessions in which you acquired imagery with camera times corresponding to our target request times: Acraman Impact Crater, Australia; Afar Rift Zone, Ethiopia (confirmed); Bosumtwi Impact Crater, Ghana; Gulf of Fonseca (confirmed), Central America; Heard Island, Southern Ocean; Honolulu, HI (confirmed); Kingman Reef & Palmyra Atoll, central Pacific (confirmed); Mississippi Delta Region (confirmed); Mount Nyiragongo, central Africa; Patagonian Glaciers (confirmed); Pilcomayo River dynamics, Argentina; Santa Barbara Coast, CA (confirmed); and Volcan Colima, Mexico (confirmed). This is an incredible response to our target requests! We have also noted with interest your views of cities at night (PAO has inquired about your very nice views of Chicago). We'd also like to acknowledge the striking views of Australia and New Zealand acquired on your time. Thank you for your energetic and enthusiastic support of our payload!'

CEO photo targets uplinked for today were Heard Island (Heard Island is a bleak, uninhabited, and mountainous island located in the Southern Ocean; about two-thirds of the way from Madagascar to Antarctica. Its mountains are covered in glaciers and dominated by Mawson Peak, a 9,006 ft high complex volcano which forms part of the Big Ben massif. A long thin spit named "Elephant Spit" extends from the east of the island. As was the case yesterday, this pass was just after midday, however weather is marginal. Looking well right of track and using the long lens for details), Goat Paddock Impact Crater (this small, young impact site is located in northwestern Australia. It lies just north of the Great Sandy Desert and near where the Margaret River breaks out of the Kimberly Plateau. Although Goat Paddock is only 5.1-km in diameter it is less than 50 million years old and is a fairly distinct feature. ISS approach was from the SW in mid-afternoon light. Looking just left of track with the long lens), and Patagonian Glaciers (ISS pass was near nadir over the southern part of the Southern Patagonian Ice Field; the larger of two icefields in the southern Andes Mountains of Chile and Argentina. The crew had several passes over Patagonia today, but the weather looked to be best for a late morning pass. As the station approached the coast from the west, the crew was to look for views and details of the less-photographed glaciers on the western and southern flanks of the ice field).

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use